how do I solve this equation and what are the steps. I can't seem to figure it out. Thank You
45/d=5

Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

45/d=5
Multiply by d
45 = 5d
Divide by 5
9 = d

Answer by stanbon(75887)

45/d=5
-----
One-way to work it:
Cross-multiply
45 = 5d
Divide by 5 to get:
d = 9
-------
Another way to work it:
Invert both sides to get "d" into a numerator:
d/45 = 1/5
Cross-multiply
d = 45(1/5)
d = 9
=================
Cheers,
Stan H.